please help me solve these problems
b/3=-3/4

z2+z=2

Answer by checkley77(12844)

b/3=-3/4
4b=-3*3
4b=-9
b=-9/4 ans.
----------------------------
z2+z=2
z^2+z-2=0
(z+2)(z-1)=0
z+2=0
z=-2 ans.
z-1=0
z=1 ans.